### v2

# Energy-Efficient Channel Decoding for Wireless Federated Learning: Convergence Analysis and Adaptive Design

Abstract
One of the most critical challenges for deploying distributed learning solutions, such as federated learning (FL), in wireless networks is the limited battery capacity of mobile clients. While it is a common belief that the major energy consumption of mobile clients comes from the uplink data transmission, this paper presents a novel finding, namely the channel decoding operation also contributes significantly to the overall energy consumption of mobile clients in FL. Motivated by this new observation, we propose an energy-efficient adaptive channel decoding scheme that leverages the intrinsic robustness of FL to model errors. In particular, the robustness is exploited to reduce the energy consumption of channel decoders at mobile clients by adaptively adjusting the number of decoding iterations. We theoretically prove that wireless FL with communication errors can converge at the same rate as the case with error-free communication as long as the bit error rate (BER) is properly constrained. An adaptive channel decoding scheme is then proposed to improve the energy efficiency of wireless FL systems. Experimental results demonstrate that the proposed method maintains the same learning accuracy while reducing the channel decoding energy consumption by 20% when compared to existing approaches.
Index Terms:
Federated learning (FL), energy efficiency, channel decoding, low-density parity checking (LDPC), bit error rate (BER).
I Introduction
I-A Motivation
Deep learning has achieved remarkable success in numerous realms, such as computer vision and natural language processing [1]. However, the conventional centralized learning paradigm encounters the data isolation issue due to the rising concern about privacy [2]. As a result, uploading the distributed data, which may contain privacy-sensitive information, for centralized learning becomes prohibited. To address this issue, federated learning (FL) was proposed [3] to enable the distributed training of deep learning models without sharing the data of users (e.g., IoT devices). However, to fully unlock the potential of FL, we need to tackle several critical challenges, such as data and device heterogeneity [4, 5, 6], as well as the communication bottleneck[7]. Another issue of wireless FL is the energy consumption of mobile clients due to heavy computation and communication workloads for training and transferring the deep neural network (DNN) parameters[8].
Many innovative approaches have been proposed to reduce the energy consumption of mobile clients in FL from different perspectives. For instance, reinforcement learning was utilized in [9] to identify near-optimal client selection and resource allocation to reduce both the convergence time and energy consumption of mobile clients. In [10], the trade-off between learning accuracy and total energy consumption of users was optimized through computational and communication resource allocation. In [11], the transmission and computation energy was minimized by bandwidth allocation and power control. Besides system-level resource management, techniques such as model sparsification [12] and quantization [13, 14, 15, 16, 17, 18] have been proposed to minimize the communication overhead, thus reducing the energy consumption of the transmitters. Authors of [19] minimized the transmission energy consumption of mobile clients while maintaining the learning performance by optimizing the transmit power, user selection and uplink resource allocation. Authors of [20] adopted quantization and periodic aggregation to reduce the transmission volume and energy, which also provided a theoretical framework about the impact of quantization and periodic aggregation on the convergence rate of FL. Model sparsification and quantization techniques are also beneficial for minimizing the processing energy consumption [21, 8]. Authors of [21] proposed to introduce sparsity at training time to reduce the on-device training energy. Authors of [8] reduced the local computation energy by quantization on both DNN weights and activations.
However, most works on energy-efficient FL focused on the energy consumption of mobile transmitters while few of them paid attention to the energy consumption of channel decoders in mobile receivers. As will be shown in Section II-B, channel decoders consume a substantial proportion of energy because of the iterative processes to decode massive DNN parameters in each communication round of wireless FL.
To address the research gap, in this work, we first quantitatively analyze the energy consumption of different components in a mobile unit and show that the channel decoder contributes significantly to the overall energy consumption. Building on this observation, an energy-efficient adaptive channel decoding scheme is proposed for wireless FL systems, by leveraging the fact that FL is robust to model errors due to its statistical learning nature[22]. Furthermore, by extending the convergence analysis in[20] to the scenario of FL with downlink communication erros, we investigate the relationship between the bit error rate (BER) and the convergence rate of wireless FL, and determine the maximum tolerable BER that can maintain the same convergence rate as the case with error-free communication. Finally, we propose a practical approach to adjust the number of channel decoding iterations in each communication round to achieve the maximum tolerable BER, while reducing the energy consumption of mobile clients in wireless FL.
I-B Contributions
This paper investigates the impact of channel decoding on the energy efficiency of wireless FL. Our major contributions are summarized as follows:
  • We theoretically analyze the relationship between BER and the convergence rate of wireless FL. Specially, an upper bound of the average expected gradient norm of the global training loss with respect to BERs over communication rounds is derived by extending the analysis in [20], based on which, the maximum tolerable BER to achieve the same convergence rate as error-free FL is determined.
  • Based on the theoretical analysis, an adaptive low-density parity checking (LDPC) decoding mechanism is proposed, which controls the maximum number of decoding iterations to satisfy the BER requirement in every communication round and maintain the target convergence rate as error-free wireless FL, while reducing the energy consumption of channel decoders in wireless FL.
  • Experimental results on both IID and non-IID datasets show that the proposed adaptive scheme effectively reduces the energy consumption of channel decoders by 20% when compared with the traditional schemes which set a fixed maximum number of channel decoding iterations, while guaranteeing the same learning performance.
I-C Organization
The remainder of this paper is organized as follows. Section II introduces the system model and presents a comprehensive energy consumption analysis for mobile units. Section III develops an energy-efficient adaptive LDPC decoding scheme for wireless FL systems by exploiting the relationship between the convergence rate and the achieved BER of LDPC decoding. Experimental results are presented in Section IV. Finally, conclusions are drawn in Section V.
II System Model and Preliminaries
In this section, we first introduce the wireless FL system. An overview regarding the energy consumption of different components at a mobile unit (i.e., client) is then presented, which highlights the significant contribution of channel decoding to the total energy consumption of wireless FL. For ease of reference, we list the key notations in Table I.
TABLE I: Key Notations
Notation Definition
$\mathbf{w}$ DNN model
$\Delta\mathbf{w}_{r}^{k}$ Local model update of client $k$ at the $r$-th communication round
$f$ The loss function of training
$f^{*}$ Minimum value of the global training loss
$R$ The total communication rounds
$E$ Number of local training steps in each communication round
$\sigma_{L}^{2}$ Variance of local gradient estimator
$\sigma_{G}^{2}$ Dissimilarity between local and global gradient
$n$ Number of total clients
$N$ Each DNN parameter is with $N$ bit representation
$d$ Dimension of the DNN model
$T$ Total training iterations, i.e., $T=RE$.
$L$ The $L$-smooth parameter
$b_{r}$ Bit error rate in the $r$-th communication round
$\eta$ Learning rate
$Q$ Maximum number of LDPC decoding iterations
$M_{r}$ Range of the global model parameters in the $r$-th communication round
$\hat{M}$ Maximum range of global model parameters for all communication rounds
(a) Wireless FL system: Multiple mobile clients are coordinated by one remote server through noisy wireless channels, where model updates and model weights are communicated in the uplink and downlink, respectively.
(b) System operations: The server performs model aggregation, digitalization, LDPC encoding, and model broadcasting. Mobile clients perform LDPC decoding, local training, and uploading the model updates.
Figure 1: Wireless FL in digital communication network and its operations.
II-A Wireless FL
Fig. 1(a) depicts a wireless FL system, where model training of different clients is coordinated by a remote server. The FL problem can be formulated as follows [23, 24, 25]:
$\min_{\mathbf{w}\in{\mathbb{R}^{d}}}f(\mathbf{w})\triangleq\frac{1}{n}\sum_{k=1}^{n}f_{k}(\mathbf{w}),$ (1)
where $\mathbf{w}\in{\mathbb{R}^{d}}$ denotes the DNN model with $d$ parameters, $n$ represents the number of mobile clients, $f_{k}\left(\cdot\right)$ is the local loss function of the $k$-th client, and $f\left(\cdot\right)$ is the global loss function.
Figure 2: Digitalization scheme for model parameters. This part gives an example, say [min, max] is digitalized by 3-bits and shows how each value inside is mapped to ”000”, ”001”, ”…”, ”111”.
To train a shared global model, the server communicates with the clients for multiple rounds through noisy wireless channels. As shown in Fig. 1(b), in each communication round, the global model at the server is broadcast to the clients. Each client then performs local training based on the received global model and transmits the model updates after local training to the server for aggregation. Specifically, in the $r$-th communication round, the global model is digitalized by $N$ bits, encoded by a channel code, and broadcast to clients. The model parameter digitalization scheme is illustrated in Fig. 2, where the range of [min, max] is divided into $12^{N}-$ 1 intervals and the width of each interval is $\Delta=\frac{\text{max-min}}{2^{N}-1}$. In this paper, we assume $N$ is large enough so the continuous and digital representations of model parameters have negligible differences. The model parameters received by the $k$-th client in the $r$-th communication round is given as follows:
$\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r},\ k=1,\cdots,n,\ r=0,\cdots,R-1,$ (2)
where $\mathbf{\tilde{w}}_{r}$ denotes the distorted version of $\mathbf{w}_{r}$ due to channel decoding errors. The impact of channel decoding errors can be quantified by considering the different error bit positions in the $N$-bit representations of model parameters, which will be analyzed in Section III-A. Each client updates the received global model by conducting $E$ steps of local mini-batch Stochastic Gradient Descent (SGD) with a learning rate $\eta$ as follows:
$\begin{split}\mathbf{w}^{k}_{r,e+1}&=\mathbf{w}^{k}_{r,e}-\eta{\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k}),\\ &e=0,\cdots,E-1,\ k=1,\cdots,n,\end{split}$ (3)
where ${\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k})$ denotes the gradient computed from a mini-batch datasets $\xi_{r,e}^{k}$ at the $k$-th client.
By completing $E$ steps of local training, the local model update of the $k$-th client can be expressed as follows:
$\begin{split}\Delta\mathbf{w}^{k}_{r}&=\mathbf{w}^{k}_{r,E}-\mathbf{w}^{k}_{r,0},\\ &r=0,\cdots,R-1,\ k=1,\cdots,n,\end{split}$ (4)
which is also digitalized, encoded, and uploaded to the server. For simplicity, we assume an orthogonal uplink channel is available for each client. After receiving these local updates, denoted as $\Delta\mathbf{\tilde{w}}^{k}_{r}$, the server performs model aggregation to update the global model as follows:
$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum_{k=1}^{n}\Delta\mathbf{\tilde{w}}^{k}_{r},\ r=0,\cdots,R-1.$ (5)
In this work, we focus on the downlink communication rather than the uplink since our objective is to design novel channel decoding schemes to save the energy consumption of mobile clients. For the uplink, the channel decoding happens at the server, where a powerful channel decoder as well as a sufficient number of decoding iterations can always be applied to achieve satisfactory BER performance. As such, we consider $\Delta\mathbf{\tilde{w}}_{r}^{k}=\Delta\mathbf{w}_{r}^{k}$ in the following.
II-B Energy Consumption Evaluation
There are two main energy-consuming operations for clients participating in wireless FL, namely model training and model communication. For the latter, it is believed that the energy is mostly consumed by the data transmission. In this section, we evaluate the energy consumption of different components, taking the training of AlexNet for image classification tasks as an example [26]. The analysis is based on actual measurements obtained from fabricated chips published in recent years [27, 28, 29]. For fair comparisons, all data are collected for chips fabricated in the 65 nm complementary metal-oxide semiconductor (CMOS) process.
In[27], a fabricated AI hardware accelerator was reported to achieve the energy efficiency of 13.7 mJ/epoch for 8-bit quantized AlexNet. With five epochs of local training, the processing energy consumption in each communication round is 68.5 mJ. In [28], a fifth-generation (5G) New Radio wireless transmitter and receiver were reported to achieve the energy efficiencies of 47.9 pJ/b and 33.3 pJ/b for transferring uncoded data, respectively, under a communication distance of around 200 meters. Therefore, the energy required to transmit and receive the 8-bit AlexNet with 60 million parameters in a single communication round is about 39 mJ. Assuming a code rate of 1/2, the communication energy consumption for coded transmission of the AlexNet model parameters in each round is 78 mJ. Besides, the LDPC decoder in [29] achieves the energy efficiency of 20.1 pJ/b/iteration, which was measured with 10 decoding iterations on average. Accordingly, the energy consumption of the LDPC decoder for decoding the AlexNet model parameters is 96.5 mJ. The detailed comparison is presented in Table II, which indicates that the energy consumption of LDPC decoding accounts for a large proportion of the overall energy consumption at clients in wireless FL.
TABLE II: Energy Consumption of Main Components at a Mobile Unit in Wireless FL for Training an AlexNet
Component Energy efficiency Energy consumption
AI Chip 13.7 mJ/epoch[27] 68.5 mJ
RF Transceiver 81.2 pJ/b[28] 78 mJ
LDPC Decoder 20.1 pJ/b/iter[29] 96.5 mJ
II-C LDPC Decoding
In this work, we use LDPC codes as an example to illustrate how adaptive channel decoding can help reduce the energy consumption of wireless FL. In LDPC, the iterative decoding process continues until $Q$ decoding iterations, denoted as the maximum number of decoding iterations, are completed [30]. Such iterative LDPC decoding processing leads to high energy consumption. To save the decoding energy, early termination techniques are usually adopted. Specifically, if the decoded codewords can fully pass the parity check, the decoding process will stop even if not all the $Q$ decoding iterations has been completed[31]. The early termination scheme can save decoding iterations without compromising the BER performance. The energy consumption of LDPC decoding depends on the target BER and thus the pre-defined maximum number of decoding iterations. As a result, to minimize the energy consumption, it is essential to determine the maximum tolerable BER by FL systems, considering its unique statistical learning nature.
III Energy-Efficient LDPC Decoding Design
In this section, we propose an energy-efficient LDPC decoding scheme to minimize the decoding energy consumption while guaranteeing the same convergence rate for wireless FL as the case with error-free communication. For that purpose, we explore the relationship between BER and convergence rate, and that between the maximum number of decoding iterations and BER. By connecting these two relationships, we demonstrate how to regulate the LDPC decoding scheme.
III-A Effect of BER on the Convergence Rate of Wireless FL
With error-free communication, the convergence rate of FL with non-convex loss functions is proved as $\mathcal{O}(1/\sqrt{T})$ [20], where $T$ is the total number of SGD iterations. In this subsection, we investigate how BER affects the convergence rate of FL and determine the maximum admissible BER for maintaining the convergence rate of $\mathcal{O}(1/\sqrt{T})$.
Unlike FL with analog communication[32, 33, 34, 35, 36, 37, 38], quantifying the impact of BER on FL with digital communication is not straightforward. In particular, we need to consider different cases of bit errors and analyze the impact of different bit errors, e.g., most significant bits and least significant bits.
To conduct the theoretical analysis, we take the following assumptions that are widely adopted in the literature.
Assumption 1.
($L$-Smoothness [13, 20]) The loss function $f\left(\cdot\right)$ is $L$-smooth, i.e., $\forall\mathbf{w},\mathbf{v}$, we have $||\nabla f(\mathbf{w})-\nabla f(\mathbf{v})||\leq L||\mathbf{w}-\mathbf{v}||$.
Assumption 2.
(Unbiased and Variance-bounded Local Gradient Estimator [39]) Let $\xi_{k}$ denote the batch of randomly sampled data at client $k$. The gradient estimator $\nabla f_{k}(\mathbf{w};\xi_{k})$ is unbiased, i.e., $\mathbb{E}_{\xi_{k}}[\nabla f_{k}(\mathbf{w};\xi_{k})]=\nabla f_{k}(\mathbf{w})$, and has bounded variance, i.e., $\mathbb{E}_{\xi_{k}}[||\nabla f_{k}(\mathbf{w};\xi_{k})-\nabla f_{k}(\mathbf{w})||^{2}]\leq\sigma_{L}^{2}$.
Assumption 3.
(Bounded Local and Global Gradient Dissimilarity [39]) The dissimilarity between the local and global gradient is bounded, i.e., $||\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})||^{2}\leq\sigma_{G}^{2}$.
Assumption 4.
(Independent Bit Errors [40]) Each bit in the bitstream of the decoded model parameters encounters error independently.
With the $N$-bit representation, each model parameter has $12^{N}-$ 1 possible distorted patterns. However, our empirical results in Fig. 4 show that it is sufficient to focus on cases with at most one bit error of each model parameter. In particular, Fig. 4 compares the actual model error with the analytical model error expression in (7) assuming at most 1-bit error, which match closely with various BER. Therefore, we impose the following assumption to simplify our analysis.
Assumption 5.
(At Most 1-bit Error of each Model Parameter) There is at most 1-bit error in the $N$-bit representation of each model parameter.
Based on Assumptions 4 and 5, we are able to establish the relationship between the decoded model at a client and the broadcasted global model in the following lemma.
Lemma 1.
Denote $\tilde{\mathbf{w}}$ as the decoded model of $\mathbf{w}$. Based on Assumptions 4 and 5, the bias of each dimension in $\mathbf{w}$ is given as
$\tilde{w}_{d}=w_{d}+\delta\left(w_{d}\right).$ (6)
The mean square error caused by bit errors is given as
$\mathbb{E}[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}]=\frac{d\left(4^{N}-1\right)}{3\left(2^{N}-1\right)^{2}}\cdot b(1-b)^{N-1}M^{2},$ (7)
where $b$ is the BER.
Proof.
Please refer to Appendix V-A. ∎
Figure 3: The difference between received noisy model and broadcasted perfect global model.
(a) BER=$10^{-2}$
(b) BER=$10^{-3}$
(c) BER=$10^{-4}$
Figure 4: Validation of the model error formula by the experiment of vanilla-CNN on Fashion-MNIST dataset. The experimental setting is in Section IV-A. In every communication round, we compute the actual model error between the noisy DNN model and perfect model, and compare with the value from the analytical function in (7). It is shown that the value from analytical function is very close to the actual value, for different BER levels and model range in different communication rounds.
After understanding the effect of BER, we analyze the convergence property of wireless FL with communication errors.
Lemma 2.
With Assumptions 1-5 and the learning rate $\eta$ satisfies
$1-\frac{n+L}{n}L\eta-2L^{2}\eta^{2}E(E-1)\geq 0,$ (8)
the average expected gradient norm of $\{\nabla f(\mathbf{\bar{w}}_{r,e})\}$ is bounded as follows:
$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}\\ &+\frac{L(L+2)d\hat{M}^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}\\ &+\frac{d\hat{M}^{2}}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n},\end{split}$ (9)
where $f^{*}$ is the minimum global training loss, $\hat{M}$ is the maximum range of DNN weights over all communication rounds, $b_{r}$ is the BER in the $r$-th communication round, and $\mathbf{w}_{0}$ is the initial global model.
Proof.
Please refer to Appendix V-B. ∎
Remark 1.
It is worthy noting that the first term on the RHS of (9) captures the distance between the initial and minimum global training loss. The second and third terms indicates how bit errors in each communication round affect learning convergence. The last two terms stem from non-IID data among clients and the random mini-batch sampling.
We next simplify (9) by applying $\eta=\frac{1}{L\sqrt{RE}}=\frac{1}{L\sqrt{T}}$, which satisfies (8) with a sufficient large value of $R$.
Theorem 1.
With Assumptions 1-5 and a sufficient large value of $R$ such that $\eta=\frac{1}{L\sqrt{RE}}$ satisfies (8), (9) can be simplified as follows:
$\begin{split}\frac{1}{T}&\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ \leq&\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}\\ &+\frac{C_{2}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}}{\sqrt{T}}+\frac{C_{3}}{\sqrt{T}}+\frac{C_{4}}{T},\end{split}$ (10)
where $C_{0}\triangleq 2L(\mathbb{E}f(\mathbf{w}_{0})-f^{*})$, $C_{1}\triangleq\frac{L^{2}(L+2)dM^{2}(4^{N}-1)}{3(2^{N}-1)^{2}}$, $C_{2}\triangleq LdM^{2}$, $C_{3}\triangleq\frac{(L+1)(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}$, $C_{4}\triangleq\frac{L^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}$.
According to Theorem 1, in order to guarantee the convergence rate of $\mathcal{O}(1/\sqrt{T})$, it is crucial to control the BER in each communication round such that the RHS of (10) reduces to zero at a rate of $\mathcal{O}(1/\sqrt{T})$. Since the explicit relationship between $b_{r}$ and $r$ is difficult to obtain, we propose a simple BER regulation method in Corollary 1. As such, wireless FL systems can achieve the same convergence rate as those with error-free communication.
Corollary 1.
If the BER in each communication round is regulated as
$b_{r}\propto\frac{1}{(r+1)^{2}},r=0,\cdots,R-1,$ (11)
the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, i.e., wireless FL converges to a local optimal solution of the global training loss funcation at a rate of $\mathcal{O}(1/\sqrt{T})$.
Proof.
When $b_{r}\propto\frac{1}{(r+1)^{2}}\triangleq\frac{\vartheta}{(r+1)^{2}}$, to prove the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, it is sufficient to verify
$\lim_{T\rightarrow\infty}\frac{\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}=0$ (12)
due to the fact that $0\leq\left(b_{r}(1-b_{r})^{N-1}\right)^{2}\leq b_{r}(1-b_{r})^{N-1}\leq 1$ for $0\leq b_{r}\leq 1$. Specifically, since
$\begin{split}0&\leq\frac{\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}\leq\frac{\sum_{r=0}^{R-1}b_{r}}{\sqrt{T}}=\vartheta\frac{\sum_{r=0}^{R-1}1/(r+1)^{2}}{\sqrt{T}}\\ &=\frac{\vartheta}{\sqrt{T}}\left(1+\sum_{r=1}^{R-1}\frac{1}{(r+1)^{2}}\right)\\ &<\frac{\vartheta}{\sqrt{T}}\left(1+\sum_{r=1}^{R-1}\frac{1}{r(r+1)}\right)=\frac{\vartheta}{\sqrt{T}}\left(2-\frac{1}{R}\right)\\ &=\frac{\vartheta}{\sqrt{T}}\left(2-\frac{E}{T}\right),\end{split}$ (13)
we conclude the proof by letting $T\rightarrow\infty$ for both sides of (13) and applying the Sandwich Theorem. ∎
Corollary 1 indicates that gradually decreasing the BER at a rate of $\mathcal{O}(1/(r+1)^{2})$ can guarantee wireless FL converges at a rate of $\mathcal{O}(1/\sqrt{T})$. The proposed BER regulation rule achieves an improving communication quality with communication rounds. This is consistent with the solution developed in [36], which improves the communication quality by increasing the signal-to-noise ratio (SNR) with communication round.
III-B BER vs. Number of Iterations
The analysis in Section III-A indicates that the convergence rate of wireless FL with error-free communication can be achieved if the BER is properly controlled. In this subsection, we analyze the relationship between BER and the maximum number of LDPC decoding iterations, based on which, we develop an adaptive decoding scheme to maintain the same convergence rate.
1 Input: The initial global model $\mathbf{w}_{0}$
2 Output: The converged global model $\mathbf{w}_{R}$
3 for  Communication round $r=0,\cdots,R-1$ do
4       for Client index $k=1,\cdots,n$ do
5             $b_{r}\propto(1/(r+1)^{2})$ with $b_{0},b_{R-1}$
6             $Q=f^{-1}(b_{r})$
7             for $decoding.iter\leq Q$ do
8                  one iteration of LDPC decoding
9                   if Early termination is satisfied then
10                        break
11                   end if
12                  $decoding.iter+=1$
13             end for
14            output: $\mathbf{\tilde{w}}_{r}$
15             for local training step $e=0,...,E-1$ do
16                  $\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r}$
17                   $\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta\nabla f_{k}(\mathbf{w}^{k}_{r,e};\xi^{k}_{r,e})$
18             end for
19            $\Delta\mathbf{w}_{r}^{k}=\mathbf{w}^{k}_{r,E}-\mathbf{w}_{r,0}^{k}$
20       end for
21      $\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum^{n}_{k=1}\Delta\mathbf{w}_{r}^{k}$
22 end for
Algorithm 1 Regulating LDPC decoder in FL
It has been shown in [41] that for message-passing-based channel decoders to achieve a fraction of 1-$\epsilon$ of the channel capacity over a binary erasure channel, the relationship between the decoding complexity (denoted as $\chi_{D}$) and BER (denoted as $b$) can be modeled as follows:
$\chi_{D}=\mathcal{O}\left(\log\frac{1}{b}\right)+\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right).$ (14)
According to [42], the relationship between the decoding complexity and the maximum number of decoding iterations ($Q$) can be modeled as follows:
$\chi_{D}=Q\cdot\chi_{0},$ (15)
where $\chi_{0}$ denotes the decoding complexity per iteration. By combining (14) and (15), it leads to the following relationship between the maximum number of decoding iterations and BER:
$Q=\frac{1}{\chi_{0}}\mathcal{O}\left(\log\frac{1}{b}\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right).$ (16)
In the next subsection, we develop an adaptive scheme to control the maximum number of decoding iterations for clients in a wireless FL system based on Corollary 1 and (16).
III-C Decoding Iteration Control
In this subsection, we show how to adaptively control the maximum number of LDPC decoding iterations over communication rounds to achieve the desired BER.
In principle, by combining (11) and (16), we obtain the following relationship between the maximum number of LDPC decoding iterations and communication rounds:
$Q_{r}=\frac{1}{\chi_{0}}\mathcal{O}\left(2\log\ r\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right),$ (17)
which provides a guideline to select the maximum number of LDPC decoding iterations in the $r$-th communication round. Specifically, $Q_{r}$ can be set as $\alpha\log r+\beta$, where $\alpha$ and $\beta$ are constants. However, (17) cannot be directly utilized because $\alpha$ and $\beta$ are code-specific and depend on the receive SNR as well as channel decoding algorithms. Thus, we propose a practical approach by the mapping relationship which captures the relationship between BER and the maximum number of decoding iterations at different levels of receive SNR. Entries in the mapping relationship can be obtained through empirical measurements specific to the LDPC code used in the wireless FL system. Fig. 5 shows the exampled simulation results for a rate 1/2 LDPC code with a block length of 1008 bits and binary phase shift keying (BPSK) modulation over an additional white Gaussian noise (AWGN) channel with $\text{SNR}=1.5$ and 2.5 dB. Data in this figure can be stored in memory [43]. From the mapping relationship, we can get the corresponding maximum number of decoding iteration for a target BER for clients with different receive SNR.
Figure 5: The relation between BER and LDPC maximum decoding iterations under different SNRs.
The wireless FL procedures with adaptive LDPC decoding are summarized in Algorithm 1. The algorithm starts with an initial target BER, denoted as $b_{0}$, which is set to a relatively high value, i.e., $10^{-1}$ in our experiments. The target BER at the last communication round, denoted as $b_{R-1}$, is set as a small value, i.e., $10^{-4}$ in our experiments. The concrete expression of $b_{r}$ in our experiment is $b_{r}=\frac{0.0999R^{2}}{(R^{2}-1)(r+1)^{2}}+\frac{10^{-4}R^{2}-0.1}{R^{2}-1}$. For each communication round, the BER is scaled according to (11), and the corresponding maximum number of LDPC decoding iterations in that round is obtained from the mapping relationship.
Remark 2.
Note that the number of entries in the mapping relationship depends on the possible discrete BER and SNR values. Since this number is not large, the memory overhead for storing the mapping relationship is expected to be small. Additionally, reading data from the memory takes constant time and does not introduce additional LDPC decoding latency.
IV Experimental Results
In this section, we validate the effectiveness of the proposed adaptive LDPC decoding algorithm via numerical experiments. First, we introduce the experimental settings. Then, the performance of the proposed algorithm, including test accuracy and energy consumption, is compared with a baseline with a fixed number of LDPC decoding iterations.
Figure 6: Fixed decoding schemes: On the client side, setting the LDPC maximum decoding iterations to 6, 12, 18, 24, and 52 can achieve a BER of around $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$, respectively.
IV-A Experimental Settings
(a) Experiment of LeNet-300-100 on MNIST.
(b) Experiment of vanilla-CNN on Fashion-MNIST.
(c) Experiment of CNN on CIFAR-10.
(d) Experiment of vanilla-CNN on non-IID Fashion-MNIST.
Figure 7: Experiments on different datasets showing how fixed BER affects FL accuracy.
(a) Consumed decoding iteration for each frame.
(b) The distributed for different consumed iterations.
Figure 8: Example of vanilla-CNN on Fashion-MNIST. The 8-bit quantized vanilla CNN model goes through AWGN channel with $E_{b}/N_{0}$ of 2.5dB, and is decoded by LDPC of block-length of 1008 bit.
System Setups: We consider an FL system with 10 clients. All clients are assumed to be selected in every communication round and they communicate with the server using rate 1/2 min-sum LDPC code with a code length of 1008 bits[44]. Besides, binary phase shift keying (BPSK) is adopted as the modulation scheme. In each communication round, the number of local SGD steps is set as 5 with the learning rate and batch size as 0.01 and 64, respectively. The receive SNR of all the wireless channels is assumed to be 2.5 dB.
Datasets: We conduct experiments on three dataset, including MNIST[45], Fashion-MNIST[46], and CIFAR-10[47]. The training datasets are distributed among clients in an IID manner. We further consider one experiment where the Fashion-MNIST data are non-IID among clients. In particular, each client was assigned with two categories of training data, similar to the setup in [24]. The test datasets are used for validation at the server.
DNN Models: For the MNIST dataset, the LeNet-300-100 model[48] is employed, which is a fully connected neural network consisting of two hidden layers and one softmax layer, with approximately 0.27 million parameters. For the Fashion-MNIST dataset, a vanilla CNN [3] with two 5×5 convolution layers, one fully connected layer, and one softmax layer is employed, comprising around 0.58 million parameters. For the CIFAR-10 dataset, a 7-layer CNN model with four 3×3 convolution layers, two fully connected layers, and one softmax layer is employed, which contains approximately 9.07 million parameters.
IV-B Learning Performance of Conventional Schemes
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
Figure 9: Experiment for LeNet-300-100 on MNIST dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve the same final accuracy of 98%, and reduce the decoding energy by 6.7mJ.
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
Figure 10: Experiment for vanilla CNN on IID Fashion-MNIST dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 91%, and reduce the decoding energy by 14.1mJ.
In this subsection, experimental results for the baseline scheme are presented. As shown in Fig. 6, it leads to approximately constant BERs over the training process.
Next, we investigate the impact of a constant BER on the performance of wireless FL. Fig. 7 summarizes the test accuracy with different maximum numbers of decoding iterations. It can be observed that for all the four FL tasks, when the BER is higher than $10^{-3}$, the learning performance was significantly impaired. Lowering the BER to $10^{-4}$ leads to improved accuracy and faster convergence towards wireless FL with error-free communication. However, further decreasing the BER to $10^{-5}$ does not yield any significant improvement in test accuracy.
IV-C Learning Performance of the Proposed Scheme
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
Figure 11: Experiment for 7-layer CNN on CIFAR-10 dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 77.6%, and reduce the decoding energy by 190mJ.
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
Figure 12: Experiment for vanilla CNN on non-IID Fashion-MNIST dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 54.1%, and reduce the decoding energy by 13.9mJ.
In this subsection, the proposed adaptive LDPC decoding scheme is compared to the baseline scheme. Specifically, the learning process of the proposed scheme is assigned with an initial BER of $10^{-1}$ and a final BER value of $10^{-4}$. The desired BER at every communication round is calculated via (11) and the maximum number of LDPC decoding iterations is determined according to the mapping with BER as stated in Section III. In the experiments, the initial and final maximum numbers of LDPC decoding iterations are thus set to 6 and 24, respectively. For the baseline scheme, a maximum of 24 LDPC decoding iterations are used in each communication round, which corresponds to a BER of $10^{-4}$.
Fig. 8(a) shows an example on the Fashion-MNIST (IID) datasets to illustrate the actual number of decoding iterations for each frame of data bits. It can observed that only a few frames exhaust all the decoding iterations which is set to a large value of 52. Note that the decoding energy consumption is proportional to the actual number of decoding iterations used for the FL tasks. From the distribution on the actual number of decoding iterations shown in Fig. 8(b), it is observed that 18 decoding iterations are sufficient for most cases, implying the significant reduction of energy consumption.
Fig. 9 summarizes the experimental results for LeNet-300-100 on the MNIST (IID) dataset. Specifically, Fig. 9(a) shows how the maximum number of LDPC decoding iterations varies with communication rounds in the proposed scheme. While the baseline scheme with a fixed number of maximum decoding iterations results in a constant BER, the proposed scheme achieves the desired BER by adopting a linearly increasing maximum number of decoding iterations. Fig. 9(b) shows the test accuracy achieved by different schemes. It is observed that setting the maximum number of decoding iterations to 6 (which corresponds to a BER of $\sim 10^{-1}$) severely affects the learning accuracy. Similarly, when the maximum number of decoding iterations is set to 12 (which corresponds to a BER of $\sim 10^{-2}$), the training process still fails to converge. By setting the maximum number of decoding iterations as 18 (which corresponds to a BER of $\sim 10^{-3}$), the training process converges, but with a noticeable loss in test accuracy compared to the case when the maximum number of decoding iterations is 24 (which corresponds to a BER of $\sim 10^{-4}$). However, further increasing the maximum number of decoding iterations to 52 does not lead to much accuracy improvement.
The proposed scheme, which linearly increases the maximum number of decoding iterations from 6 to 24, achieves the same test accuracy as those with fixed 24 or 52 decoding iterations. Fig. 9(c) illustrates how the average number of decoding iterations over all DNN parameters varies during the training process. The proposed scheme can reduce the average number of decoding iterations and thus the overall energy consumption. Fig. 9(d) depicts the BER of the proposed scheme over communication rounds, which scales as $\mathcal{O}(1/(r+1)^{2})$ from $10^{-1}$ to $10^{-4}$. Fig. 9(e) shows the final test accuracy of different schemes at the 50-th communication round. It can be observed that the proposed scheme achieves the same accuracy as the baseline schemes that always maintain a very low BER of $10^{-4}$ and $10^{-5}$ throughout the training process. Fig. 9(f) illustrates the decoding energy consumption on clients under different schemes. It is obtained by multiplying the total number of LDPC decoding iterations with the LDPC decoding energy efficiency reported in [29]. Compared to the baseline scheme with a maximum of 24 decoding iterations, the proposed scheme reduces the overall decoding energy consumption by 19.4%.
TABLE III: Experimental Results of Different Schemes. Our scheme can achieve similar accuracy while saving around 20% of the decoding energy compared to the best-fixed scheme.
Final learning accuracy
Fixed-18 Fixed-24 Fixed-52 Ours
MNIST 96.74% 97.96% 98.06% 98.02%
Fashion-MNIST 86.18% 90.63% 91.24% 91.16%
CIFAR-10 76.72% 77.38% 77.57% 77.68%
non-IID Fashion 51.3% 54.1% 54.3% 54.2%
Total consumed decoding energy (mJ)
Fixed-18 Fixed-24 Fixed-52 Ours
MNIST 33.87 34.51 34.68 27.81
Fashion-MNIST 74.23 76.10 76.29 62.17
CIFAR-10 925.10 948.44 951.36 757.80
non-IID Fashion 74.22 76.11 76.27 62.38
Figs. 10, 11, and 12 present the experimental results for Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets, respectively. The results are similar to those on the MNIST (IID) dataset in Fig. 9, showing that adapting the maximum number of LDPC decoding iterations can achieve similar or even slightly better accuracy compared to the baseline scheme. This observation was also reported in [49, 50, 51], which suggests that adding annealed noise to the DNN model could enhance its generalization ability. Moreover, the proposed scheme reduces the LDPC decoding energy consumption on the Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets by 18.3%, 20.1%, and 18.0%, respectively. Table III provides a summary of the experimental data.
V Conclusion
In this paper, we considered the energy consumption issue in wireless federated learning (FL) over digital communication networks. Our study recognized that channel decoding contributes to a large portion of the energy consumption at mobile clients in wireless FL systems. To achieve energy-efficient wireless FL, we first show that by properly controlling the bit error rate (BER) over different communication rounds, wireless FL can achieve the same convergence rate as FL with error-free communication. Based on this analysis, we proposed an energy-efficient LDPC decoding scheme to regulate the maximum number of decoding iterations over communication rounds. Experimental results demonstrated that the proposed method can reduce approximately 20% of the total energy consumed by LDPC decoding, while maintaining the same or even a slightly higher learning accuracy compared to the scheme without adaptive decoding iteration control.
Appendix
V-A Proof of Lemma 1
We express a decoded model $\tilde{\mathbf{w}}$ in terms of $\mathbf{w}$ as follows:
$\tilde{\mathbf{w}}=\mathbf{w}+\bm{\delta}\left(\mathbf{w}\right),$ (18)
where $\bm{\delta}\left(\mathbf{w}\right)$ is a $d$-dimensional function of $\mathbf{w}$ denoting the model distortion caused by bit errors. Under Assumption 4, each entry of $\bm{\delta}\left(\mathbf{w}\right)$ depends only on $w_{d}$, i.e., the $d$-th dimension of $\mathbf{w}$. Denote $\delta\left(w_{d}\right)$ as the $d$-th dimension of $\bm{\delta}\left(\mathbf{w}\right)$. Therefore, the $d$-th dimension of $\tilde{\mathbf{w}}$ is given as
$\tilde{w}_{d}=w_{d}+\delta\left(w_{d}\right).$ (19)
Under Assumptions 4 and 5, the 1-bit error may occur at different bits, leading to different levels of model distortion. Let $w_{d}^{(i)}\in\{0,1\}$ be the $i$-th digit ($i=0$ and $N-1$ respectively denote the least and most significant bits) in the $N$-bit representation of model parameter $w_{d}$. Hence,
$\mathbb{E}\left[\tilde{w}_{d}|w_{d}\right]=w_{d}+b\left(1-b\right)^{N-1}\Delta\sum_{i=0}^{N-1}\left(1-2w_{d}^{(i)}\right)2^{i},$ (20)
The expected bias of $w_{d}$ is bounded as follows:
$\begin{split}|\mathbb{E}\left[\tilde{w}_{d}|w_{d}\right]-w_{d}|&\leq b\left(1-b\right)^{N-1}\left(2^{N}-1\right)\Delta\\ &=b\left(1-b\right)^{N-1}M.\end{split}$ (21)
The mean square error of $w_{d}$ is given as follows:
$\begin{split}\mathbb{E}\left[||\tilde{w}_{d}-w_{d}||^{2}|w_{d}\right]&=b\left(1-b\right)^{N-1}\sum_{i=0}^{N-1}\left(\Delta 2^{i}\right)^{2}\\ &=b\left(1-b\right)^{N-1}\frac{4^{N}-1}{3(2^{N}-1)^{2}}M^{2}.\end{split}$ (22)
Therefore,
$\begin{split}\mathbb{E}\left[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}\right]&=\sum_{j=1}^{d}\mathbb{E}\left[||\tilde{w}_{j}-w_{j}||^{2}|\mathbf{w}\right]\\ &=\frac{d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\left(1-b\right)^{N}M^{2}.\end{split}$ (23)
V-B Proof of Lemma 2
We define
$\mathbf{w}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\mathbf{\tilde{w}}_{r}),$ (24)
$\mathbf{\bar{w}}_{r,e}$ $=\frac{1}{n}\sum_{k=1}^{n}\mathbf{w}_{r,e}^{k},$ (25)
for each communication round $r=0,...,R-1$ and each local iteration $e=0,\cdots,E-1$, where $\mathbf{w}_{r+1}$ denotes the updated global model and $\mathbf{\bar{w}}_{r,e}$ is the average of client models after the $e$-th local iteration.
To prove the Lemma 2, we first present Lemmas 26$\sim$5.
Lemma 3.
For $\mathbf{\bar{w}}_{r,E}$ and $\mathbf{w}_{r+1}$ defined in (24) and (25), we have
$\mathbf{w}_{r+1}-\mathbf{\bar{w}}_{r,E}=\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r},\ r=0,\cdots,R-1.$ (26)
Proof.
The proof can be obtained by substituting the definition of $\mathbf{\bar{w}}_{r,E}$ to the LHS of (26). ∎
Lemma 26 translates the difference between $\mathbf{w}_{r+1}$ and $\mathbf{\bar{w}}_{r,E}$ to the model distortion incurred by downlink bit errors.
Lemma 4.
With Assumption 1, for $r=0,...,R-1$,
$\begin{split}\mathbb{E}f(\mathbf{w}_{r+1})&\leq\mathbb{E}f(\mathbf{\bar{w}}_{r,E})+\frac{L}{2}\mathbb{E}||\mathbf{w}_{r+1}-\mathbf{\bar{w}}_{r,E}||^{2}\\ &-\mathbb{E}\left\langle\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle.\end{split}$ (27)
Proof.
The proof can be obtained using the $L$-smoothness of $f(\cdot)$, which is omitted for brevity. ∎
Lemma 27 provides an upper bound on the expected value of the global loss function at $\mathbf{w}_{r+1}$.
Lemma 5.
With Assumptions 1 - 3, for $r=0,\cdots,R-1$,
$\begin{split}&\mathbb{E}f(\mathbf{\bar{w}}_{r,E})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L}{2}\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &-\frac{\eta}{2n}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}\\ &+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (28)
Proof.
The proof can be obtained following similar lines for Lemma 7 in [20] by further considering the non-IID data distribution among clients (i.e., Assumption 3), which is omitted for brevity. ∎
Lemma 28 provides an upper bound on the expected value of the global loss function at $\bar{\mathbf{w}}_{r,E}$ in terms of the expected model distortion caused by downlink bit errors, i.e., $\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}$. By applying Lemmas 26 and 28 to Lemma 27, we obtain the following inequality:
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+L\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &-\frac{\eta}{2n}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (29)
By leveraging $2\left\langle\mathbf{a},\mathbf{b}\right\rangle=\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}-\|\mathbf{a}-\mathbf{b}\|^{2}\leq\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}$ holds for any $\mathbf{a}$, $\mathbf{b}$, an upper bound for the third term on the RHS of (29) is derived as follows:
$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{\mathbb{E}\|\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E})\|^{2}}{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{a})}{\leq}\frac{L^{2}}{2}\mathbb{E}\|\mathbf{\bar{w}}_{r,E}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\mathbf{w}_{r})\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\tilde{\mathbf{w}}_{r})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{=}\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{\eta}{n}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\nabla f_{k}(\mathbf{w}_{e}^{k};\xi_{r,e}^{k})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\end{split}$ (30)
where (a) is due to the $L$-smoothness of $f(\cdot)$ and (b) is due to the local updating rules. Based on the Assumptions 2 and 3, we have
$\begin{split}\mathbb{E}\|\nabla&f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\|^{2}=\mathbb{E}\|\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\\ &-\nabla f_{k}(\mathbf{w})+\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})+\nabla f(\mathbf{w})\|^{2}\\ &\leq\sigma_{L}^{2}+\sigma_{G}^{2}+\mathbb{E}\|\nabla f(\mathbf{w})\|^{2}.\end{split}$ (31)
By combing (31) and (30), the bound in (30) can be further updated as
$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{L^{2}\eta^{2}}{2n^{2}}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\|^{2}+\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{\leq}\frac{L^{2}\eta^{2}}{2n^{2}}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}+\frac{L^{2}\eta^{2}E(\sigma_{G}^{2}+\sigma_{L}^{2})}{2n}\\ &\ \ \ +\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2},\end{split}$ (32)
Substituting (32) in (29), we have
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L(L+2)}{2}\mathbb{E}\|\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &-\frac{\eta}{2n}\left[1-\frac{n+L}{n}L\eta-2L^{2}\eta^{2}E(E-1)\right]\\ &\cdot\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (33)
For any $\eta$ that satisfies (8), we have
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L(L+2)}{2}\mathbb{E}\|\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r}\|^{2}\\ &+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (34)
Summing (34) over $r=0,\cdots,R-1$, rearranging the terms, and multiplying $\frac{2}{\eta RE}$ on both sides, we obtain the following expression:
$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-\mathbb{E}f(\mathbf{w}_{R}))}{\eta RE}\\ &+\frac{L(L+2)}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}+\frac{1}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.\end{split}$ (35)
Based on (23) and (21) respectively, we have
$\begin{split}\mathbb{E}\left[||\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}||^{2}\right]&=\mathbb{E}\left[\mathbb{E}\left[||\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}||^{2}|\mathbf{w}_{r}\right]\right]\\ &=\frac{d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b_{r}\left(1-b_{r}\right)^{N-1}\cdot M_{r}^{2},\end{split}$ (36)
$\begin{split}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}&=\sum_{j=1}^{d}\mathbb{E}\left[|\delta(w_{r,d})|^{2}|{w}_{r,d}\right]\\ &\leq\sum_{j=1}^{d}\mathbb{E}\left[(b_{r}(1-b_{r})^{N-1}M)^{2}|{w}_{r,d}\right]\\ &\leq d(b_{r}(1-b_{r})^{N-1})^{2}M_{r}^{2}.\end{split}$ (37)
By substituting (36) and (37) into (35), and denote $f^{*}$ as the minimum global training loss, we have
$\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}$
$+\frac{L(L+2)d(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}M_{r}^{2}$
$+\frac{d}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}M_{r}^{2}$
$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.$ (38)
Denoting $\hat{M}=max(M_{0},M_{1},\cdots,M_{R-1})$, we have
$\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}$
$+\frac{L(L+2)d\hat{M}^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}$
$+\frac{d\hat{M}^{2}}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}$
$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.$ (39)